What is the Hardly-Weinberg equation and what do the element represent?
1
Expert's answer
2019-07-30T04:51:05-0400

The Hardy-Weinberg equation is the fundamental concept in population genetics (the study of genetics in a defined group). It is a mathematical equation describing the distribution and expression of alleles (forms of a gene) in a population, and it expresses the conditions under which allele frequencies are expected to change.

According to this equation:

p = the frequency of the dominant allele (represented here by A)

q = the frequency of the recessive allele (represented here by a)

For a population in genetic equilibrium:

p + q = 1.0 (The sum of the frequencies of both alleles is 100%.)

(p + q)2 = 1

so

p2 + 2pq + q2 = 1

The three terms of this binomial expansion indicate the frequencies of the three genotypes:

p2 = frequency of AA (homozygous dominant)

2pq = frequency of Aa (heterozygous)

q2 = frequency of aa (homozygous recessive)